# Safety Mirage: How Spurious Correlations Undermine VLM Safety Fine-tuning

### v1

@@ 4 The Risks of Spurious Correlations in VLM Safety Fine-Tuning
Spurious text features and spurious correlations. As discussed in Sec. 3, the effectiveness of current VLM safety fine-tuning methods heavily relies on the curation of high-quality, dual-modality safety datasets. As a result, the safety capabilities of fine-tuned VLMs (i.e., their abilities to prevent harmful content generation) are primarily learned from safety labels (i.e., safe responses) introduced in the fine-tuning datasets. For example, in VLGuard [23], safety labels for unsafe text queries and/or unsafe images are assigned as rejection responses, such as “I’m sorry, I cannot assist with that request…” against the unsafe text query in Figs. 1-2. Additionally, safety labels may correspond to standard VLM responses when processing safe text-image inputs.
At first glance, the use of safety labels appears appropriate. However, a hidden bias may arise when these safety labels become strongly correlated with spurious features in the input data, particularly within textual queries, which are the focus of this work. Here the term “spurious features” refers to non-essential textual words in inputs that do not contribute to the fundamental meaning or task-relevant aspects of the input query. For example, in Fig. 2, the word “What” or “Share” at the beginning of the input query can be considered a spurious feature because it does not directly relate to the query’s actual content and can be easily substituted with other question words. In contrast, “core” features (such as “crime”) are more informative, representing content-related words that capture the true intent and meaning of the query. Therefore, we define spurious correlations as the (unexpected) strong associations between spurious input features and the assigned labels in the safety fine-tuning dataset.
The above conceptualization of spurious correlations is inspired by conventional analyses in image classification [69, 70, 71], where spurious features correspond to background or non-object pixels, while core features directly relate to the target object. In these studies, spurious correlations arise when the target label is linked to spurious features, rather than solely to core features. In the context of VLMs, while several studies have explored spurious correlations [72, 73], their focus has been primarily on spurious visual features within the image modality. To the best of our knowledge, the spurious correlations caused by textual features in VLM safety fine-tuning have not been investigated before.
In this work, we identify two types of spurious correlations in VLM safety fine-tuning. (a) Non-rejection bias: Certain words (like “What” in Fig. 2(a)) in text queries become spuriously correlated with non-rejection responses. As a result, incorporating these words into an original query can easily jailbreak fine-tuned VLMs. (b) Rejection bias: Certain words (like “Share” in Fig. 2(b)) in text queries become spuriously correlated with rejection responses, causing the fine-tuned VLM to exhibit over-prudence.
Refer to caption Refer to caption
(a) Top words in safe queries (b) Top words in unsafe queries
@@
Figure 3: Frequency of question-initiating words in VLGuard queries: (a) words in safe queries associated with non-rejection responses, (b) words in unsafe queries associated with rejections.
To determine the spurious textual features and the spurious correlations (a)-(b), we analyze the frequency of words used in train-time text queries that lead to non-rejection responses (i.e., generating textual content in response to safe queries in the training set) and rejection responses (i.e., predefined refusal answers against unsafe queries in the training set), respectively. Fig. 3 presents the most frequently occurring starting words in text queries from the VLGuard dataset, categorized based on their tendency to elicit non-rejection responses or rejection responses. As shown, the question word “what” predominantly correlates with non-rejection responses, appearing in over 80% of safe queries where the model generates a response. In contrast, the question-initiating words “can” and “share” in unsafe queries are strongly associated with rejection responses, with over 50% of its occurrences leading to a rejection. Compared to “can”, which appears in both safe and unsafe queries, the word “share” is only used in unsafe queries to elicit rejection.
@@
One-word jailbreaking. Recognizing the non-rejection bias (i.e., the spurious correlation between certain querying words like “what” and non-rejection responses) as shown in Fig. 3, an adversary can exploit this spurious correlation to jailbreak safety fine-tuned VLMs. Formally, let $q$ denote the original unsafe text query that VLM avoids generating unsafe content for. And let $q^{\prime}=\mathrm{w}_{\mathrm{adv}}+q$ denote the jailbreaking attack, where $\mathrm{w}_{\mathrm{adv}}$ is the adversarial perturbation chosen based on the non-rejection bias-inducing querying word (e.g., “what” in Figs. 1, 2, and 3), and the operation $+$ signifies either word insertion as a prefix to $q$ or a simple starting word replacement in $q$. We refer to the above simple attack strategy as the one-word jailbreaking attack.
@@
Figure 4: ASR of $K$-shot one-word attack for varying $K$, evaluated before and after applying “What”-initialized one-word attack to jailbreak the safety fine-tuned VLM (LLaVA-v1.5-7B-Mixed [23]).
[…]
One-word over-prudence. Similar to the one-word jailbreaking attack, when a rejection bias-inducing word (like “Share” in Fig. 3) is introduced into text queries, we observe that even benign modified queries can still prevent the fine-tuned VLM from generating any meaningful output. To amplify the over-prudence rate, we can also apply the multi-shot strategy (i.e., “Share”-based one-word modification to different paraphrased versions of the benign input query). The results for the multi-shot setting are presented in Appendix A. Notably, even in the one-shot setting, the “Share”-initialized benign query modification already achieves a 90% over-rejection rate against safe text-image queries.
@@ 5 Enhancing VLM Safety through Unlearning
To mitigate spurious correlations embedded in the safety fine-tuning dataset, one potential solution, motivated by the analysis in Sec. 4, is to eliminate the dependence of safety fine-tuning on safety labels (e.g., rejection responses to unsafe queries). This necessitates shifting from supervised fine-tuning to a label-free, unsupervised setting for safety alignment. Machine unlearning (MU) [27, 28, 29] provides an ideal solution in this context, as it is designed to remove the undesired influence of harmful data or knowledge from a pre-trained model while preserving its normal utility.
Although unlearning has been applied to VLM safety fine-tuning in prior work [17, 30, 31], its unique advantage and application in circumventing spurious correlations within fine-tuning datasets remains unexplored. To this end, we adapt two state-of-the-art MU approaches, originally developed for large language models (LLMs)–representation misdirection unlearning (RMU) [26] and negative preference optimization (NPO) [25]–to the context of VLM unlearning.
The proposed VLM unlearning follows the generic formulation of (2), but with the following key modifications. First, the fine-tuning loss over the unsafe dataset $\mathcal{D}_{\mathrm{u}}$ is replaced with an unlearning objective $\ell_{\mathrm{u}}$ that relies solely on the unsafe data features (text-image queries) in $\mathcal{D}_{\mathrm{u}}$, without depending on the safety labels. In our work, we define the unlearning loss $\ell_{\mathrm{u}}$ based on the principles of RMU and NPO, respectively. The RMU-based unlearning objective aims to map the intermediate features of unsafe data $x\in\mathcal{D}_{\mathrm{u}}$ (to be forgotten) to random features. This ensures that the model no longer retains meaningful representations of the unsafe data. The objective is given by
@@
where $M_{{\bm{\theta}}}(\cdot)$ represents certain intermediate-layer representations of ${\bm{\theta}}$, $c$ is a hyperparameter that controls activation scaling , and $\mathbf{v}$ is a random vector drawn from a standard uniform distribution. We remark that, unlike RMU for LLM unlearning [26], we carefully adjusts the representation layer selection and tunes the hyperparameter $c$ to better suit the unlearning process in VLMs.
@@
Table 1: Experiment results evaluating safety, over-prudence, and utility of safety fine-tuned VLMs. Safety is quantified by ASR (attack success rate) on unsafe input queries, evaluated before and after a 3-shot one-word attack (i.e., “what”-based prefix in Fig. 1) that promotes non-rejection bias. Over-prudence is measured by RR (rejection rate) on safe input queries, evaluated before and after using 1-shot, one-word over-prudence modification (i.e., “share”-based prefix in Fig. 1). Here, “Before” and “After” denote the performance prior to and following the respective one-word modification. Utility is assessed by the accuracy (Acc.) on four VQA benchmarks: VQAv2, TextVQA, ScienceQA, and VizWiz. Results are presented for models under full fine-tuning and LoRA fine-tuning settings, with safety fine-tuning approaches including Unsafe-Filter, Mixed-SFT, Posthoc-SFT, NPO-Unlearning, and RMU-Unlearning.
@@
In addition to RMU, we also employ NPO [25] to model the unlearning objective $\ell_{\mathrm{u}}$, which treats unsafe data designated for unlearning as “negative” examples in a direct preference optimization framework [76]. The NPO-based unlearning loss is then given by
$\begin{array}[]{c}\ell_{\mathrm{u}}({\bm{\theta}};\mathcal{D}_{\mathrm{u}})=\mathbb{E}_{\mathbf{x}\in\mathcal{D}_{\mathrm{u}}}\left[-\frac{2}{\beta}\log\sigma\left(-\beta\log\left(\frac{\pi_{{\bm{\theta}}}(\mathbf{x})}{\pi_{\mathrm{ref}}(\mathbf{x})}\right)\right)\right],\end{array}$ (5)
where $\sigma(\cdot)$ the sigmoid function, $\beta>0$ is the temperature parameter , $\pi_{{\bm{\theta}}}$ denotes the prediction probability of the model ${\bm{\theta}}$ given the unsafe input $\mathbf{x}$, and $\pi_{\mathrm{ref}}$ represents the reference model given by the initial model prior to unlearning. The rationale behind NPO is to fine-tune the VLM ${\bm{\theta}}$ to force it to deviate from the reference model when processing unsafe inputs. Following (2), VLM unlearning also requires the retain loss $\ell_{\mathrm{r}}$ to preserve the model utility in normal tasks. This is achieved using the standard VLM training loss over the safe text and safe image data in the fine-tuning set.
As demonstrated in Sec. 6, compared to conventional supervised safety fine-tuning, unlearning-based approaches produce safer responses by generating irrelevant or neutral outputs rather than merely rejecting unsafe queries. Moreover, unlearning reduces the model’s reliance on rejection responses, alleviating over-prudence and enabling more appropriate handling of benign inputs.
6 Experiments
6.1 Experiment Setups
Datasets and models.
We consider four VLM safety datasets: VLGuard [23], MM-SafetyBench [14], SPA-VL [21], and Figstep [15]. To assess the utility of safety fine-tuned VLMs, we also conduct evaluations on representative visual question-answering (VQA) datasets, including VQAv2 [77], TextVQA [78], VizWiz [79], and ScienceQA [80]. For model selection, we adopt LLaVA-v1.5-7B and LLaVA-v1.5-13B [6, 7] as our primary VLMs.
Safety fine-tuning setups and baselines. In our experiments, we choose VLGuard as the training dataset for VLM safety fine-tuning. When implementing the MU-based fine-tuning approaches, i.e., RMU in (3) and NPO in (5), we use the unsafe input-output pairs from VLGuard as the unsafe dataset ($\mathcal{D}_{\mathrm{u}}$) to be unlearned. Here we employ Llama-2-13B-Chat to confirm the harmfulness of the original model’s responses against the unsafe input queries. Additionally, the safe query-answer pairs from VLGuard are used to construct the retain dataset ($\mathcal{D}_{\mathrm{r}}$) in (2). For additional details, we refer readers to Appendix  B.1-B.2.
Besides MU-based VLM safety fine-tuning, we include a series of popular supervised safety fine-tuning approaches as baselines. (1) Mixed-SFT [23]: Supervised fine-tuning (SFT) using a mixed fine-tuning strategy on VLGuard. (2) Posthoc-SFT [23, 64]: SFT using a post-hoc fine-tuning approach on VLGuard. (3) Unsafe-Filter: SFT performed on a clean training set, where LLaMA-Guard-3-11B-Vision [81] is used to filter unsafe data from pre-training datasets.
Evaluation setups. For a fine-tuned VLM, its safety performance is primarily evaluated using the attack success rate (ASR), defined as the proportion of unsafe responses generated for unsafe input queries over the total number of queries. The unsafe input queries are considered in two scenarios. (1) Before attack: The original unsafe input queries from VLM safety datasets including VLGuard, MM-SafetyBench, SPA-VL, and Figstep. (2) After attack: We apply the proposed one-word attack to the original unsafe queries. By default, our one-word jailbreaking attack is implemented using a 3-shot approach, as illustrated in Fig. 4. Specifically, three paraphrased versions of the text query (while preserving the adversarial perturbation) are posed for each image. The attack is considered successful if any of the generated responses is unsafe. To validate the harmfulness of the responses generated by the fine-tuned model, we employ the multi-modal Qwen2.5-VL-7B-Instruct as a judge. The details of the judge prompt can be found in Appendix  B.4.
To investigate the over-prudence issue in safety fine-tuned VLMs, we measure the rejection rate (RR), which quantifies the proportion of benign test-time queries that are unnecessarily rejected by the fine-tuned model. These benign queries consist of both safe text queries and safe images provided by the VLM safety datasets. Similar to ASR, RR is evaluated both before and after exploiting spurious correlations to encode rejection bias. Furthermore, we evaluate the general utility of safety fine-tuned VLMs by measuring their prediction accuracy (Acc) on standard downstream VQA tasks, including VQAv2, TextVQA, ScienceQA, VizWiz.

### v2

@@ 4 The Risks of Spurious Correlations in VLM Safety Fine-Tuning
Spurious text features and spurious correlations. As discussed in Sec. 3, current VLM safety fine-tuning methods heavily depend on high-quality, dual-modality safety datasets. Consequently, the safety capabilities of fine-tuned VLMs (i.e., their abilities to prevent harmful content generation) are primarily learned from safety labels (i.e., safe responses) introduced in the fine-tuning datasets. For example, in VLGuard, unsafe queries are assigned as rejection responses, such as “I’m sorry, I cannot …”. Similarly, SPA-VL also selects rejection-based answers for unsafe queries.
Refer to caption Refer to caption Refer to caption
(a) Top words in VLGuard safe   (b) Top words in VLGuard unsafe (c) Top words in SPA-VL
@@
Figure 3: Frequency of question-initiating words across training corpora: (a) VLGuard safe queries with non-rejection responses, (b) VLGuard unsafe queries with rejection responses, (c) SPA-VL queries.
At first glance, the use of safety labels appears appropriate. However, hidden bias may arise when safety labels become strongly correlated with spurious features in the input data, particularly within textual queries, which are the focus of this work. Here the term “spurious features” refers to non-essential features in inputs (primarily for texts in this work) that do not contribute to the fundamental meaning or task-relevance of the input query, in contrast to the “core” features. For example, in Fig. 1, the initial words “What” or “Share” function as spurious features, since they are not directly related to the query’s actual content and can be easily substituted with other question words. In contrast, core features (such as “crime”) are more informative, representing content-related words that capture the true meaning of the query. Therefore, we define spurious correlations as the (unexpected) strong associations between spurious input features and the safety labels in the fine-tuning dataset. The above conceptualization of spurious correlations is inspired by and remains consistent with conventional spurious correlation analyses in image classification (Sagawa et al., 2020), where background pixels serve as spurious features while object pixels represent core content.
In this work, we identify two types of spurious correlations in VLM safety fine-tuning. (a) Non-rejection bias: Certain words (like “What” in Fig. 1(b)) in text queries become spuriously correlated with non-rejection responses. So, incorporating these words into an original query can easily jailbreak fine-tuned VLMs. (b) Rejection bias: Certain words (like “Share” in Fig. 1(b)) in text queries become spuriously correlated with rejection responses, causing the fine-tuned VLM to exhibit over-prudence.
To reveal these biases, we analyze the frequency of question-initiating words used in training queries that lead to non-rejection responses (i.e., generating contents for safe queries) and rejection responses (i.e., predefined refusal answers for unsafe queries). Fig. 3 presents the most frequent starting words across VLGuard and SPA-VL. In VLGuard, the question word “what” predominantly correlates with non-rejection responses, appearing in over 80% of safe queries where the model generates a response. By contrast, words like “can” and “share” in unsafe queries are closely tied to rejection responses, with “share” exclusively occurring in unsafe contexts and leading to rejection in more than half of its cases. For SPA-VL, we again observe a strong dominance of queries beginning with “what.” Unlike VLGuard safe queries, SPA-VL often pairs such queries with both rejection and non-rejection responses. Preference optimization strengthens alignment with the safety-labeled rejections, yet “what”-initiated queries remain inherently associated with unsafe, non-rejection responses.
@@
One-word jailbreaking. Recognizing the non-rejection bias (i.e., the spurious correlation between certain words like “what” and non-rejection responses) as shown in Fig. 3, an adversary can exploit this bias to jailbreak safety fine-tuned VLMs. Formally, let $q$ be an unsafe text query that the VLM would normally refuse. We construct a paraphrased version $q^{\prime}$ to rewrite $q$ such that it begins with a chosen adversarial word $\mathrm{w}_{\mathrm{adv}}$ (e.g., “what” in Figs. 1, 2, and 3). We refer to this strategy, rephrasing unsafe queries to enforce a biased starting word, as the one-word jailbreaking attack.
In practice, we find that repeatedly applying the one-word attack, by integrating $\mathrm{w}_{\mathrm{adv}}$ with paraphrased versions of the original input query $q$ (up to $K$ times), can significantly improve the attack success rate (ASR). We refer to this strategy as $K$-shot one-word attack. Fig. 4 shows the ASR of the $K$-shot one-word attack on the safety fine-tuned VLM. Even $K=1$ yields 29% ASR, compared to near 0% for the original unsafe queries in VLGuard. ASR exceeds 50% for $K\geq 3$ and approaches 90% as $K$ increases. In contrast, paraphrased queries without the “What”-trigger remain ineffective, confirming that the bias-inducing word is essential to the attack’s success. The effectiveness of the one-word attack can be interpreted through the lens of backdoor attacks (Gao et al., 2020; Saha et al., 2020), where the bias-inducing word “What” functions as a trigger that shortcuts text queries to non-rejection responses during safety fine-tuning. Additional results on LLaVA-v1.5-7B-PPO-30K are provided in Appendix B.1.
@@
Figure 4: ASR of $K$-shot one-word attack for varying $K$, evaluated before and after applying the “What”-initialized one-word attack to jailbreak the safety fine-tuned VLM (LLaVA-v1.5-7B-Mixed (Zong et al., 2024)) on VLGuard unsafe.
Figure 5: $R$ vs. $K$-shot one-word modification, evaluated before and after applying the “Share” initialized one-word modification to safe input queries, causes the over-prudence phenomenon of the safety fine-tuned VLM LLaVA-v1.5-7B-Mixed model (Zong et al., 2024) on VLGuard safe.
One-word over-prudence. Mirroring the jailbreaking case, inserting a rejection-bias word (e.g., “Share”, Fig. 3) induces over-prudence in fine-tuned VLMs: even benign queries prefixed with this word often yield no meaningful output. A multi-shot procedure, where a benign query is rewritten into several “Share”-initialized variants, further amplifies the effect (Fig. 5). Notably, even a single modification already triggers a 90% over-rejection rate on safe text–image queries, highlighting how dataset-induced starting-word biases can severely degrade utility. Additional results for LLaVA-v1.5-7B-PPO-30K in Appendix B.2.
@@ 5 Enhancing VLM Safety through Unlearning
To mitigate spurious correlations in the safety fine-tuning dataset (Sec. 4), a natural solution is to eliminate reliance on safety labels, thereby necessitating shifts to a label-free setting for safety alignment. Machine unlearning (MU) (Liu et al., 2025; Cao and Yang, 2015; Bourtoule et al., 2021) provides an ideal solution in this context, as it is designed to remove the undesired influence of harmful data or knowledge from a pre-trained model while preserving its normal utility.
Although unlearning has been applied to VLM in prior work (Chen et al., 2025; Huo et al., 2025), its unique advantage in addressing spurious correlations remains underexplored. Therefore, we adapt two state-of-the-art MU approaches from LLMs, representation misdirection unlearning (RMU) (Li et al., 2024a) and negative preference optimization (NPO) (Zhang et al., 2024b), to VLM safety field.
The proposed VLM unlearning follows the generic formulation of (2), but with the following key modifications. First, the fine-tuning loss on the unsafe dataset $\mathcal{D}_{\mathrm{u}}$ is replaced with an unlearning objective $\ell_{\mathrm{u}}$ that relies solely on the unsafe data features (text-image queries) in $\mathcal{D}_{\mathrm{u}}$, without depending on the safety labels. In our work, we define the unlearning loss $\ell_{\mathrm{u}}$ based on the principles of RMU and NPO, respectively. The RMU-based unlearning objective aims to map the intermediate features of unsafe data $x\in\mathcal{D}_{\mathrm{u}}$ (to be forgotten) to random features. This ensures that the model no longer retains meaningful representations of the unsafe data. The objective is given by
@@
where $M_{{\bm{\theta}}}(\cdot)$ represents certain intermediate-layer representations of $\bm{\theta}$}, $c$ is a hyperparameter that controls activation scaling , and $\mathbf{v}$ is a random vector drawn from a standard uniform distribution. We remark that, unlike RMU for LLM unlearning (Li et al., 2024a), we carefully adjusts the representation layer selection and tunes the hyperparameter $c$ to better suit the unlearning process in VLMs.
In addition to RMU, we also employ NPO (Zhang et al., 2024b) to model the unlearning objective $\ell_{\mathrm{u}}$, which treats unsafe data designated for unlearning as “negative” examples in a direct preference optimization framework (Rafailov et al., 2023). The NPO-based unlearning loss is then given by
$\begin{array}[]{c}\hskip-8.53581pt\ell_{\mathrm{u}}({\bm{\theta}};\mathcal{D}_{\mathrm{u}})=\mathbb{E}_{\mathbf{x}\in\mathcal{D}_{\mathrm{u}}}\left[-\frac{2}{\beta}\log\sigma\left(-\beta\log\left(\frac{\pi_{{\bm{\theta}}}(\mathbf{x})}{\pi_{\mathrm{ref}}(\mathbf{x})}\right)\right)\right],\end{array}$ (5)
where $\sigma(\cdot)$ the sigmoid function, $\beta>0$ is the temperature parameter , $\pi_{{\bm{\theta}}}$ denotes the prediction probability of the model $\bm{\theta}$} given the unsafe input $\mathbf{x}$, and $\pi_{\mathrm{ref}}$ represents the reference model given by the initial model prior to unlearning. The rationale behind NPO is to fine-tune the VLM $\bm{\theta}$} to force it to deviate from the reference model when processing unsafe inputs.
In addition, unlearning-based safety fine-tuning on VLMs requires the retain loss $\ell_{\mathrm{r}}$ in (2) to preserve utility on normal tasks. Unlike LLM unlearning, which relies solely on MU-specific retain objectives, we find that directly applying MU objectives to VLMs often leads to instability or even model collapse. To address this, we design $\ell_{\mathrm{r}}$ as a composite of two terms:
$\ell_{\mathrm{r}}({\bm{\theta}};\mathcal{D}_{\mathrm{r}})=\ell_{\mathrm{ft}}({\bm{\theta}};\mathcal{D}_{\mathrm{r}})+\alpha\ell_{\mathrm{mu},\mathrm{r}}({\bm{\theta}};\mathcal{D}_{\mathrm{r}}),$ (6)
where $\ell_{\mathrm{ft}}$ denotes the standard fine-tuning loss of the base VLM, which stabilizes optimization and ensures consistent training dynamics, while $\ell_{\mathrm{mu},\mathrm{r}}$ is the MU-specific retain loss that enforces utility preservation (e.g., the representation loss in RMU on the utility set). Further details on $\ell_{\mathrm{r}}$ and MU implementation in the safety context are provided in Appendix C. As will be shown in Sec. 6, compared to conventional supervised safety fine-tuning, our approach yields robust safety results.
6 Experiments
6.1 Experiment Setups
Datasets and models.
We consider four VLM safety datasets: VLGuard (Zong et al., 2024), MM-SafetyBench (Liu et al., 2024c), SPA-VL (Zhang et al., 2024a), and Figstep (Gong et al., 2023). To assess the utility of safety fine-tuned VLMs, we also conduct evaluations on representative visual question-answering (VQA) datasets, including VQAv2 (Goyal et al., 2017), TextVQA (Singh et al., 2019), VizWiz (Gurari et al., 2018), and ScienceQA (Lu et al., 2022). For model selection, we adopt LLaVA-v1.5-7B and LLaVA-v1.5-13B (Liu et al., 2023, 2024a) as our primary VLMs.
Safety fine-tuning setups and baselines. In our experiments, we use VLGuard as the training dataset for VLM safety fine-tuning, where its training split is divided into two parts: unsafe and safe. The safe query–answer pairs form the retain dataset ($\mathcal{D}_{\mathrm{r}}$) used for computing $\ell_{\mathrm{r}}$ in (6). For the unlearning loss $\ell_{\mathrm{u}}$, the construction of the unsafe dataset differs by method. For NPO (5), we only use the unsafe queries as $\mathcal{D}_{\mathrm{u}}$. In contrast, for RMU (3), we concatenate each unsafe query with harmful responses selected by Llama-2-13B-Chat, pairing them to form the input to be unlearned. Additional implementation details are provided in Appendix C & D.1.
Besides MU-based VLM safety fine-tuning, we include a series of popular supervised safety fine-tuning approaches as baselines. (1) Mixed-SFT (Zong et al., 2024): Supervised fine-tuning (SFT) using a mixed fine-tuning strategy on VLGuard. (2) Posthoc-SFT (Zong et al., 2024; Taori et al., 2023): SFT using a post-hoc fine-tuning approach on VLGuard. (3) Unsafe-Filter: An SFT baseline where unsafe samples are excluded by using LLaMA-Guard-3-11B-Vision (Chi et al., 2024) to filter unsafe text–image pairs from LLaVA’s pre-training, refer to Appendix  D.2.
Evaluation setups. We assess safety performance using the attack success rate (ASR). For each unsafe query, model outputs are classified into three categories: (1) rejection responses, (2) non-rejection but irrelevant responses, and (3) unsafe responses directly related to the unsafe query. ASR is computed in two steps. First, we measure the rejection rate across all responses. For the remaining non-rejection outputs, we use the multi-modal Qwen2.5-VL-7B-Instruct model as a judge to determine whether a response is related to the unsafe query. Any related response is counted as unsafe. The full judging prompt is provided in Appendix D.3. We report ASR under two scenarios: (1) Before attack: using the original unsafe queries from safety benchmarks. (2) After attack: applying our proposed one-word jailbreaking attack to these unsafe queries. By default, the attack is implemented in a 3-shot setting, where three paraphrased versions of each query (with the adversarial perturbation preserved) are posed for each image, as illustrated in Fig. 4. When multiple responses ($K>1$) are generated, the attack is considered successful if any of them is unsafe.
To further analyze the over-prudence phenomenon in safety fine-tuned VLMs, we measure the rejection rate (RR), which quantifies the proportion of benign test-time queries that are unnecessarily rejected. Similar to ASR, RR is evaluated both before and after applying 1-shot one-word modifications designed to induce rejection bias, like Fig. 5. We also evaluate the general utility by measuring prediction accuracy (Acc) on standard downstream VQA benchmarks.
@@ 6.1 Experiment Setups
Table 1: Experiment results evaluating safety, over-prudence, and utility of safety fine-tuned VLMs. Safety is quantified by ASR (attack success rate) on unsafe input queries, evaluated before and after a 3-shot one-word attack (i.e., “what”-based prefix in Fig. 1) that promotes non-rejection bias. Over-prudence is measured by $R$ (rejection rate) on safe input queries, evaluated before and after using 1-shot, one-word modification (i.e., “share”-based prefix in Fig. 1). Here, “Before” and “After” denote the performance prior to and following the respective one-word modification. Utility is assessed by the accuracy (Acc.) on VQA benchmarks. Results are presented for models under full fine-tuning and LoRA fine-tuning settings, with safety fine-tuning approaches including Unsafe-Filter, Mixed-SFT, Posthoc-SFT, NPO-Unlearning, and RMU-Unlearning.